OpenStudy (anonymous):

Elena joined a golf club two years ago. She pays an annual membership fee of $925 and a greens fee of $38 each time she plays a game of golf. The function below can be used to calculate her total yearly golfing fee, f (g), in dollars. f (g) = 925 + 38g , where g represents the number of times she plays golf during the year. Last year she paid $1,913 as a total golfing fee. For how many games did she pay a greens fee? answer only please

OpenStudy (anonymous):

since f(g)=925+38g and f(g)= 1913 then 925+38g = 1913 38g = 988 g = 36

OpenStudy (anonymous):

26